# Acyclic Conjunctive Regular Path Queries are no Harder than Corresponding Conjunctive Queries

### v1

@@ 1. Introduction
We study acyclic Conjunctive Regular Path Queries (CRPQs) in the data complexity setting, i.e., where the query is fixed. We provide an output-sensitive algorithm for evaluating acyclic CRPQs, i.e., an algorithm whose complexity depends on both the input size and output size. We show that our algorithm improves upon prior algorithms for acyclic CRPQs (Abo Khamis et al., 2025; Abo Khamis et al., 2025). We also show that our complexity matches the best known output-sensitive complexity for the corresponding acyclic conjunctive queries (Hu, 2025). Hence, it is not possible to improve upon our complexity for acyclic CRPQswithout a new breakthrough on output-sensitive evaluation of acyclic conjunctive queries.
@@
Recent work (Abo Khamis et al., 2025) has introduced an output-sensitive algorithm for evaluating acyclic CRPQs. At a high-level, you can think of their algorithm as “lifting” the Yannakakis algorithm (Yannakakis, 1981) from acyclic conjunctive queries (CQs) to acyclic CRPQs. Let’s take the query $Q_{1}$ above. We cannot obtain an output-sensitive algorithm for $Q_{1}$ by simply evaluating each RPQ $\accentset{\rightsquigarrow}{R_{i}}(X,Y_{i})$ independently using the algorithm from (Abo Khamis et al., 2025) and then joining their outputs. This is because the output of each RPQ $\accentset{\rightsquigarrow}{R_{i}}(X,Y_{i})$ could be much larger than the final output of $Q_{1}$ itself. Instead, the main insight from (Abo Khamis et al., 2025) is to start by “calibrating” the RPQs $\accentset{\rightsquigarrow}{R_{i}}(X,Y_{i})$ with each other before invoking the algorithm from (Abo Khamis et al., 2025) on them and joining their outputs. The calibration filters out vertices that don’t contribute to the final output, similar to the calibration passes in the original Yannakakis algorithm (Yannakakis, 1981). Using this insight, the authors from (Abo Khamis et al., 2025) show how to evaluate any acyclic CRPQ that is free-connex in time $O(N+N\cdot\textsf{OUT}^{1/2}+\textsf{OUT})$, where OUT is the output size. This applies to the example query $Q_{1}$ above. However, acyclic CRPQs that are not free-connex turn out to be more challenging. Such queries are transformed in (Abo Khamis et al., 2025) into free-connex queries by “promoting” some non-free variables to become free. For example, $Q_{2}$ above can be transformed into a free-connex query by promoting $X$ to become free, thus obtaining $Q_{1}$. Using this approach, the authors of (Abo Khamis et al., 2025) obtain a “weak” output-sensitive algorithm for $Q_{2}$ whose runtime is $O(N+N\cdot\textsf{OUT}_{1}^{1/2}+\textsf{OUT}_{1})$ where $\textsf{OUT}_{1}$ is the output size of the transformed query $Q_{1}$, which could be much larger than the original output size $\textsf{OUT}_{2}$ of $Q_{2}$. To rewrite the runtime in terms of the original output size $\textsf{OUT}_{2}$, the authors use a weak upper bound of $\textsf{OUT}_{1}\leq\textsf{OUT}_{2}\cdot|V|$, which gives a runtime of $O(N+N\cdot\textsf{OUT}_{2}^{1/2}\cdot|V|^{1/2}+\textsf{OUT}_{2}\cdot|V|)$. This is unsatisfactory because it introduces a new parameter $|V|$ into the runtime, is still a loose bound for many input instances, and doesn’t address non-free-connex queries at a fundamental level. Instead, it over-approximates them using blackbox algorithms that were not designed for them. In the worst case, $|V|$ could be as large as $N$, making the runtime $O(N+N^{3/2}\cdot\textsf{OUT}_{2}^{1/2}+N\cdot\textsf{OUT}_{2})$, which can be significantly improved as we explain below.
@@
The above complexity collapses back to the complexity from (Abo Khamis et al., 2025) for RPQs and to the complexity from (Abo Khamis et al., 2025) for free-connex acyclic CRPQs. For non-free-connex acyclic CRPQs, our complexity is significantly better than the one from (Abo Khamis et al., 2025). For example, our complexity for $Q_{2}$ becomes $O(N+N\cdot\textsf{OUT}_{2}^{2/3}+\textsf{OUT}_{2})$, where $\textsf{OUT}_{2}$ is the output size of $Q_{2}$. This a more natural runtime expression and is also always [^6] To compare the two runtimes, note that for $Q_{2}$, we always have $\textsf{OUT}_{2}\leq N^{3}$. upper bounded by the runtime from (Abo Khamis et al., 2025), which was $O(N+N^{3/2}\cdot\textsf{OUT}_{2}^{1/2}+N\cdot\textsf{OUT}_{2})$. See Appendix C for a detailed comparison with (Abo Khamis et al., 2025).
@@ 2.1. Languages and Graphs
Given a regular expression $\accentset{\rightsquigarrow}{R}$ over an alphabet $\Sigma$, let $\Sigma^{-1}$ be a set of fresh symbols: one distinct fresh symbol $\sigma^{-1}$ for each symbol $\sigma\in\Sigma$. The inverse of $\accentset{\rightsquigarrow}{R}$, denoted $\accentset{\rightsquigarrow}{R}^{-1}$, is a regular expression over $\Sigma^{-1}$ that defines the reverse language $L(\accentset{\rightsquigarrow}{R})$ but where every symbol $\sigma$ is replaced by $\sigma^{-1}$. In particular, $\accentset{\rightsquigarrow}{R}^{-1}$ is defined inductively as follows:
@@ 2.2. Queries
A Conjunctive Query (CQ) has the form $Q(\bm{F})=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{n}(\bm{X}_{n})$, where $\textsf{vars}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\bigcup_{i\in[n]}\bm{X}_{i}$ is the set of variables, $\textsf{free}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\bm{F}\subseteq\textsf{vars}(Q)$ is the set of free variables, $\textsf{bound}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\textsf{vars}(Q)\setminus\textsf{free}(Q)$ is the set of bound variables, and $\textsf{atoms}(Q)$ is the set of atoms $R_{i}(\bm{X}_{i})$. We follow the standard semantics for CQs. The query graph of a CQ $Q$ is a hypergraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}:=\textsf{vars}(Q)$ and $\mathcal{E}:=\{\bm{X}_{i}\mid R_{i}(\bm{X}_{i})\in\textsf{atoms}(Q)\}$. We call a CQ $Q$ acyclic if its query graph is $\alpha$-acyclic.
@@
A Conjunctive Regular Path Query (CRPQ) $Q$ over an alphabet $\Sigma$ is of the form $Q(\bm{F})=\accentset{\rightsquigarrow}{R}_{1}(X_{1},Y_{1})\wedge\cdots\wedge\accentset{\rightsquigarrow}{R}_{n}(X_{n},Y_{n})$, where: each $\accentset{\rightsquigarrow}{R}_{i}$ is a regular path query; each $\accentset{\rightsquigarrow}{R}_{i}(X_{i},Y_{i})$ is an atom; $\textsf{vars}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\bigcup_{i\in[n]}\{X_{i},Y_{i}\}$ is the set of variables; $\textsf{free}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\bm{F}\subseteq\textsf{vars}(Q)$ is the set of free variables; $\textsf{bound}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\textsf{vars}(Q)\setminus\textsf{free}(Q)$ is the set of bound variables; and $\textsf{atoms}(Q)$ is the set of atoms. The answer to the CRPQ $Q$ on an input edge-labeled graph $G=(V,E,\Sigma)$ is defined as follows: A mapping $\mu:\textsf{free}(Q)\rightarrow V$ is in the answer of $Q$ if it can be extended to a mapping $\mu^{\prime}:\textsf{free}(Q)\cup\textsf{bound}(Q)\rightarrow V$ such that $G$ has a path from $\mu^{\prime}(X_{i})$ to $\mu^{\prime}(Y_{i})$ labeled by a string from $L(\accentset{\rightsquigarrow}{R}_{i})$, for every $i\in[n]$. We represent a mapping $\mu$ by the tuple $(\mu(X))_{X\in\textsf{free}(Q)}$, assuming a fixed order on the free variables.
@@
In order to define the query graph of a CRPQ, we first recall the definition of undirected multigraphs. An undirected multigraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a finite set of vertices $\mathcal{V}$ and a multiset of edges $\mathcal{E}$ such that each edge is a subset of $\mathcal{V}$ of size two or one. A edge of size one is called a self-loop. The degree of a vertex $v$ in $\mathcal{G}$ is the number of edges incident to $v$. A vertex that has degree one or zero is called a leaf. Given two vertices $v,u\in\mathcal{V}$, a path from $v$ to $u$ in $\mathcal{G}$ is a sequence of vertices $w_{0}:=v,w_{1},\ldots,w_{k}:=u$ for some $k\geq 1$ such that for every $i\in[k]$, there is an edge $\{w_{i-1},w_{i}\}\in\mathcal{E}$.
@@ 2.3. Width Measures
For any acyclic CRPQ (or acyclic CQ) $Q$, the followings hold:
@@ 3. Results
We now prove our Theorem 3.1. To that end, we will break it down into two steps: In the first step, we reduce the problem of answering an acyclic CRPQ to answering “free-leaf” CRPQs (Lemma 3.4). In the second step, we show how to answer free-leaf CRPQs efficiently (Lemma 3.5).
@@
Figure 1. An example demonstrating the reduction in Lemma 3.4. Top-left: The input acyclic CRPQ $Q$ from Eq. (11), where free variables are shaded. Top-right: The bound connected components of $Q$ (Definition 3.6). Bottom-right: The corresponding free-leaf CRPQs obtained by applying Proposition 3.10 to each component. Each free-leaf CRPQ is solved using the algorithm from Lemma 3.5, and the answers are stitched together into an acyclic full conjunctive query over only the free variables of $Q$. This query is depicted at the bottom-left, and it can be solved using Yannakakis algorithm.:
@@
(b) The query $Q$ rotated to have $D$ as the root.:
Figure 2. The free-leaf acyclic CRPQ $Q$ from Example 3.16. Free variables are shaded.:
@@
We assume for now that we already know the output size OUT of $Q$ upfront. We will show at the end of the example how to get rid of this assumption. Let $\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\textsf{OUT}^{3/4}$. (In general, $\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\textsf{OUT}^{1-1/|\textsf{free}(Q)|}$.) Moreover, let $\Delta^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\Delta+1$. For each query $P_{i}(X,\bm{Y})$, where $X$ is the first variable and $\bm{Y}$ is the set of remaining variables, we compute an $(X,\Delta^{\prime})$-restriction $\overline{P}_{i}(X,\bm{Y})$ of $P_{i}(X,\bm{Y})$ using one of Props. 3.133.14, or 3.15, in time $O(N\cdot\Delta^{\prime})$. At the very end, we compute a $(D,\Delta^{\prime})$-restriction $\overline{Q}(D,A,B,C)$ of $Q(D,A,B,C)$. A vertex $d$ is called light if $\deg_{\overline{Q}}(A,B,C|D=d)\leq\Delta$, and heavy otherwise (i.e., if $\deg_{\overline{Q}}(A,B,C|D=d)=\Delta^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\Delta+1$). The relation $\overline{Q}$ already contains all output tuples $(d,a,b,c)$ where $d$ is light. Let $H_{D}$ be the set of all heavy vertices $d$. Our next target is to report output tuples $(d,a,b,c)$ where $d\in H_{D}$. Note that because each heavy vertex $d$ in $H_{D}$ has at least $\Delta$ associated tuples $(a,b,c)$ where $(d,a,b,c)$ is in the output, the number of heavy vertices, $|H_{D}|$, is at most $\textsf{OUT}/\Delta=\textsf{OUT}^{1/4}$. (In general, we get $|H_{D}|\leq\textsf{OUT}^{1/|\textsf{free}(Q)|}$.) Using Prop. 2.14, we can assume that the query $\accentset{\rightsquigarrow}{R_{6}}(D,Z)\wedge H_{D}(Z)$ is equivalent to an RPQ $\accentset{\rightsquigarrow}{R_{6}^{\prime}}(D,Z)$. Before we continue, we replace $\accentset{\rightsquigarrow}{R_{6}}(D,Z)$ with $\accentset{\rightsquigarrow}{R_{6}^{\prime}}(D,Z)$, and from now on, we can assume that in the remaining query, the number of different $D$-values is at most $\textsf{OUT}^{1/4}$.

### v2

@@ 1. Introduction
Graph databases have become increasingly popular in recent years, with applications in various domains such as knowledge graphs, social networks, and the semantic web. These graph databases are often queried differently from traditional relational databases, using graph query languages such as SPARQL (Harris and Seaborne, 2013) and Cypher (Francis et al., 2018). Recursion is at the heart of graph query languages, and it is often expressed using regular expressions over edge labels, leading to the notion of Regular Path Queries (RPQs) and their conjunctive extensions, Conjunctive Regular Path Queries (CRPQs). CRPQs are part of SPARQL, the W3C standard for querying RDF data, including commonly used knowledge bases such as DBpedia and Wikidata (Figueira et al., 2020; Martens and Trautner, 2019). Recent analytical studies of SPARQL query logs indicate that the use of RPQs and CRPQs is on the rise (Bonifati et al., 2017). For example, RPQs currently make up 38% of unique queries to Wikidata (Bonifati et al., 2019; Malyshev et al., 2018).
In this paper, we study acyclic CRPQs in the data complexity setting, i.e., where the query is fixed. We provide an output-sensitive algorithm for evaluating acyclic CRPQs, i.e., an algorithm whose complexity depends on both the input size and output size. We show that our algorithm improves upon prior algorithms for acyclic CRPQs (Abo Khamis et al., 2025; Abo Khamis et al., 2025). We also show that our complexity matches the best known output-sensitive complexity for the corresponding acyclic conjunctive queries (Hu, 2025). Hence, it is not possible to improve upon our complexity for acyclic CRPQs  without a new breakthrough on output-sensitive evaluation of acyclic conjunctive queries.
@@
Recent work (Abo Khamis et al., 2025) has introduced an output-sensitive algorithm for evaluating acyclic CRPQs. At a high-level, you can think of their algorithm as “lifting” the Yannakakis algorithm (Yannakakis, 1981) from acyclic conjunctive queries (CQs) to acyclic CRPQs. Let’s take the query $Q_{1}$ above. We cannot obtain an output-sensitive algorithm for $Q_{1}$ by simply evaluating each RPQ $\accentset{\rightsquigarrow}{R_{i}}(X,Y_{i})$ independently using the algorithm from (Abo Khamis et al., 2025) and then joining their outputs. This is because the output of each RPQ $\accentset{\rightsquigarrow}{R_{i}}(X,Y_{i})$ could be much larger than the final output of $Q_{1}$ itself. Instead, the main insight from (Abo Khamis et al., 2025) is to start by “calibrating” the RPQs $\accentset{\rightsquigarrow}{R_{i}}(X,Y_{i})$ with each other before invoking the algorithm from (Abo Khamis et al., 2025) on them and joining their outputs. The calibration filters out vertices that don’t contribute to the final output, similar to the calibration passes in the original Yannakakis algorithm (Yannakakis, 1981). Using this insight, the authors from (Abo Khamis et al., 2025) show how to evaluate any acyclic CRPQ that is free-connex in time $O(N+N\cdot\textsf{OUT}^{1/2}+\textsf{OUT})$, where OUT is the output size. This applies to the example query $Q_{1}$ above.
Extending the approach from (Abo Khamis et al., 2025) from free-connex CRPQs to acyclic CRPQs that are not free-connex turns out to be more challenging. Such non-free-connex queries are transformed in (Abo Khamis et al., 2025) into free-connex ones by “promoting” some non-free variables to become free. For example, $Q_{2}$ above can be transformed into a free-connex query by promoting $X$ to become free, thus obtaining $Q_{1}$. Using this approach, the authors of (Abo Khamis et al., 2025) obtain a “weak” output-sensitive algorithm for $Q_{2}$ whose runtime is $O(N+N\cdot\textsf{OUT}_{1}^{1/2}+\textsf{OUT}_{1})$ where $\textsf{OUT}_{1}$ is the output size of the transformed query $Q_{1}$, which could be much larger than the original output size $\textsf{OUT}_{2}$ of $Q_{2}$. To rewrite the runtime in terms of the original output size $\textsf{OUT}_{2}$, the authors use a weak upper bound of $\textsf{OUT}_{1}\leq\textsf{OUT}_{2}\cdot|V|$, which gives a runtime of $O(N+N\cdot\textsf{OUT}_{2}^{1/2}\cdot|V|^{1/2}+\textsf{OUT}_{2}\cdot|V|)$. This is unsatisfactory because it introduces a new parameter $|V|$ into the runtime, is still a loose bound for many input instances, and doesn’t address non-free-connex queries at a fundamental level. Instead, it over-approximates them using blackbox algorithms that were not designed for them. In the worst case, $|V|$ could be as large as $N$, making the runtime $O(N+N^{3/2}\cdot\textsf{OUT}_{2}^{1/2}+N\cdot\textsf{OUT}_{2})$, which can be significantly improved as we explain below.
@@
The above complexity collapses back to the complexity from (Abo Khamis et al., 2025) for RPQs and to the complexity from (Abo Khamis et al., 2025) for free-connex acyclic CRPQs. For non-free-connex acyclic CRPQs, our complexity is significantly better than the one from (Abo Khamis et al., 2025). For example, our complexity for $Q_{2}$ becomes $O(N+N\cdot\textsf{OUT}_{2}^{2/3}+\textsf{OUT}_{2})$, where $\textsf{OUT}_{2}$ is the output size of $Q_{2}$. This is a more natural runtime expression and is also always [^6] To compare the two runtimes, note that for $Q_{2}$, we always have $\textsf{OUT}_{2}\leq N^{3}$. upper bounded by the runtime from (Abo Khamis et al., 2025), which was $O(N+N^{3/2}\cdot\textsf{OUT}_{2}^{1/2}+N\cdot\textsf{OUT}_{2})$. See Appendix C for a detailed comparison with (Abo Khamis et al., 2025).
@@ 2.1. Languages and Graphs
Given a regular expression $\accentset{\rightsquigarrow}{R}$ over an alphabet $\Sigma$, let $\Sigma^{-1}$ be a set of fresh symbols: one distinct fresh symbol $\sigma^{-1}$ for each symbol $\sigma\in\Sigma$. The inverse of $\accentset{\rightsquigarrow}{R}$, denoted $\accentset{\rightsquigarrow}{R}^{-1}$, is a regular expression over $\Sigma^{-1}$ that defines the reverse language of $L(\accentset{\rightsquigarrow}{R})$ but where every symbol $\sigma$ is replaced by $\sigma^{-1}$. In particular, $\accentset{\rightsquigarrow}{R}^{-1}$ is defined inductively as follows:
@@ 2.2. Queries
A Conjunctive Query (CQ) has the form $Q(\bm{F})=R_{1}(\bm{X}_{1})\wedge\cdots\wedge R_{n}(\bm{X}_{n})$, where $\textsf{vars}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\bigcup_{i\in[n]}\bm{X}_{i}$ is the set of variables, $\textsf{free}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\bm{F}\subseteq\textsf{vars}(Q)$ is the set of free variables, $\textsf{bound}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\textsf{vars}(Q)\setminus\textsf{free}(Q)$ is the set of bound variables, and $\textsf{atoms}(Q)$ is the set of atoms $R_{i}(\bm{X}_{i})$. We follow the standard semantics for CQs. The query graph of a CQ $Q$ is a hypergraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}:=\textsf{vars}(Q)$ and $\mathcal{E}:=\{\bm{X}_{i}\mid R_{i}(\bm{X}_{i})\in\textsf{atoms}(Q)\}$. We call a CQ $Q$ acyclic if its query graph is $\alpha$-acyclic, which means that we can construct a tree whose nodes are the sets $\bm{X}_{i}$ for $i\in[n]$ where each variable $X\in\bm{V}$ appears in a connected subtree.
@@
A Conjunctive Regular Path Query (CRPQ) $Q$ over an alphabet $\Sigma$ is of the form $Q(\bm{F})=\accentset{\rightsquigarrow}{R}_{1}(X_{1},Y_{1})\wedge\cdots\wedge\accentset{\rightsquigarrow}{R}_{n}(X_{n},Y_{n})$, where: each $\accentset{\rightsquigarrow}{R}_{i}$ is a regular expression over $\Sigma$; each $\accentset{\rightsquigarrow}{R}_{i}(X_{i},Y_{i})$ is an atom; $\textsf{vars}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\bigcup_{i\in[n]}\{X_{i},Y_{i}\}$ is the set of variables; $\textsf{free}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\bm{F}\subseteq\textsf{vars}(Q)$ is the set of free variables; $\textsf{bound}(Q)\stackrel{{\scriptstyle\text{def}}}{{=}}\textsf{vars}(Q)\setminus\textsf{free}(Q)$ is the set of bound variables; and $\textsf{atoms}(Q)$ is the set of atoms. The answer to the CRPQ $Q$ on an input edge-labeled graph $G=(V,E,\Sigma)$ is defined as follows: A mapping $\mu:\textsf{free}(Q)\rightarrow V$ is in the answer of $Q$ if it can be extended to a mapping $\mu^{\prime}:\textsf{free}(Q)\cup\textsf{bound}(Q)\rightarrow V$ such that $G$ has a path from $\mu^{\prime}(X_{i})$ to $\mu^{\prime}(Y_{i})$ labeled by a string from $L(\accentset{\rightsquigarrow}{R}_{i})$, for every $i\in[n]$. We represent a mapping $\mu$ by the tuple $(\mu(X))_{X\in\textsf{free}(Q)}$, assuming a fixed order on the free variables.
@@
In order to define the query graph of a CRPQ, we first recall the definition of undirected multigraphs. An undirected multigraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a finite set of vertices $\mathcal{V}$ and a multiset of edges $\mathcal{E}$ such that each edge is a subset of $\mathcal{V}$ of size two or one. An edge of size one is called a self-loop. The degree of a vertex $v$ in $\mathcal{G}$ is the number of edges incident to $v$. A vertex that has degree one or zero is called a leaf. Given two vertices $v,u\in\mathcal{V}$, a path from $v$ to $u$ in $\mathcal{G}$ is a sequence of vertices $w_{0}:=v,w_{1},\ldots,w_{k}:=u$ for some $k\geq 1$ such that for every $i\in[k]$, there is an edge $\{w_{i-1},w_{i}\}\in\mathcal{E}$.
@@ 2.3. Width Measures
For any acyclic CRPQ (or acyclic CQ) $Q$, the following claims hold:
@@ 3. Results
We now prove our Theorem 3.1. To that end, we will break it down into two steps: In the first step, we reduce the problem of answering an acyclic CRPQ to answering “free-leafCRPQs (Lemma 3.4). In the second step, we show how to answer free-leaf CRPQs efficiently (Lemma 3.5).
@@
Figure 1. An example demonstrating the reduction in Lemma 3.4. Top-left: The input acyclic CRPQ $Q$ from Eq. (11), where free variables are shaded. Top-right: The bound connected components of $Q$ (Definition 3.6). Bottom-right: The corresponding free-leaf CRPQs obtained by applying Proposition 3.10 to each component. Each free-leaf CRPQ is solved using the algorithm from Lemma 3.5, and the answers are stitched together into an acyclic full conjunctive query over only the free variables of $Q$. This query is depicted at the bottom-left, and it can be solved using Yannakakis algorithm.
[…]
@@
(b) The query $Q$ rotated to have $D$ as the root.
The query \$Q\$ rotated to have \$D\$ as the root.
Figure 2. The free-leaf acyclic CRPQ $Q$ from Example 3.16. Free variables are shaded.
[…]
@@
We assume for now that we already know the output size OUT of $Q$ upfront. We will show at the end of the example how to get rid of this assumption. Let $\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\textsf{OUT}^{3/4}$. (In general, $\Delta\stackrel{{\scriptstyle\text{def}}}{{=}}\textsf{OUT}^{1-1/|\textsf{free}(Q)|}$.) Moreover, let $\Delta^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\Delta+1$. For each query $P_{i}(X,\bm{Y})$, where $X$ is the first variable and $\bm{Y}$ is the set of remaining variables, we compute an $(X,\Delta^{\prime})$-restriction $\overline{P}_{i}(X,\bm{Y})$ of $P_{i}(X,\bm{Y})$ using one of Props. 3.133.14, or 3.15, in time $O(N\cdot\Delta^{\prime})$. At the very end, we compute a $(D,\Delta^{\prime})$-restriction $\overline{Q}(D,A,B,C)$ of $Q(D,A,B,C)$. A vertex $d$ is called light if $\deg_{\overline{Q}}(A,B,C|D=d)\leq\Delta$, and heavy otherwise (i.e., if $\deg_{\overline{Q}}(A,B,C|D=d)=\Delta^{\prime}\stackrel{{\scriptstyle\text{def}}}{{=}}\Delta+1$; recall that $\deg_{\overline{Q}}(A,B,C|D=d)$ cannot exceed $\Delta^{\prime}$ by definition of $(D,\Delta^{\prime})$-restriction). The relation $\overline{Q}$ already contains all output tuples $(d,a,b,c)$ where $d$ is light. Let $H_{D}$ be the set of all heavy vertices $d$. Our next target is to report output tuples $(d,a,b,c)$ where $d\in H_{D}$. Note that because each heavy vertex $d$ in $H_{D}$ has at least $\Delta$ associated tuples $(a,b,c)$ where $(d,a,b,c)$ is in the output, the number of heavy vertices, $|H_{D}|$, is at most $\textsf{OUT}/\Delta=\textsf{OUT}^{1/4}$. (In general, we get $|H_{D}|\leq\textsf{OUT}^{1/|\textsf{free}(Q)|}$.) Using Prop. 2.14, we can assume that the query $\accentset{\rightsquigarrow}{R_{6}}(D,Z)\wedge H_{D}(D)$ is equivalent to an RPQ $\accentset{\rightsquigarrow}{R_{6}^{\prime}}(D,Z)$. Before we continue, we replace $\accentset{\rightsquigarrow}{R_{6}}(D,Z)$ with $\accentset{\rightsquigarrow}{R_{6}^{\prime}}(D,Z)$, and from now on, we can assume that in the remaining query, the number of different $D$-values is at most $\textsf{OUT}^{1/4}$.